# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v6

@@ Abstract
Swarms of autonomous interactive drones can provide compelling sensing capabilities in Smart City applications, such as traffic monitoring. This paper focuses on the task assignment problem for large-scale spatio-temporal sensing by a drone swarm. However, existing approaches have distinct challenges: evolutionary distributed optimization, such as collective learning, lacks long-term adaptability in dynamic environments, while deep reinforcement learning (DRL) is limited to scale effectively due to the curse of dimensionality. Therefore, this paper proposes a novel synergetic optimization approach by integrating long-term DRL and short-term collective learning. Through this approach, each drone independently and proactively determines its flying direction and recharging location using DRL, while evolving their navigation and sensing policies through collective learning based on a structured tree communication model. Extensive experiments with datasets generated from realistic urban mobility demonstrate an outstanding performance of the proposed solution in complex scenarios. New insights show that this approach provides a win-win synthesis of short-term and long-term strategies for drone-based traffic monitoring, with short-term methods addressing training complexity and energy management, while long-term methods preserving high sensing performance.
Index Terms:
drones, UAV, navigation, sensing, recharging, distributed optimization, deep reinforcement learning
I Introduction
@@
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities [1]. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as monitoring traffic congestion [2] and mapping natural disasters [3]. Despite these advantages, drones are constrained by limited battery capacity, which impacts their spatio-temporal coverage. This limitation necessitates sophisticated autonomous control systems and precise task assignment within drone swarms. This involves the assignment of different sensing tasks to each drone while meeting the sensing requirements, drone capabilities and constraints [4]. Recently, scholars have solved the NP-hard combinatorial optimization problem of task assignment for spatio-temporal sensing, using evolutionary algorithms ranging from genetic algorithms to swarm intelligence methods [4, 5]. To enhance robustness and minimize the effects of individual drone failures, distributed optimization approaches have been introduced. These approaches allow drones to autonomously self-organize and self-assign tasks in a decentralized manner, preserving the autonomy of individual drones. However, these distributed approaches are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically.
@@
Strategic foresight in navigation and sensing is crucial for optimizing drone operations in dynamic environments, significantly influenced by flying directions and future sensor data requirements. For example, drones aware of an expected increase in traffic can proactively fly to those areas, improving tasks such as vehicle detection, even if the short-term benefits are not evident. This "slower is faster" effect emphasizes the importance of long-term planning for overall effectiveness in drone sensing. To address long-term optimization, deep reinforcement learning (DRL) has been introduced, utilizing the Bellman equation to consider cumulative rewards [6]. DRL based on deep neural networks support drones, as heterogeneous agents, to manage complex state spaces and adapt to changing environments.
@@
Therefore, in this paper, we address a novel task assignment problem for large-scale spatio-temporal sensing, focusing on the navigation (determining flight path), sensing (scheduling data collection) and recharging (choosing charging stations as destinations) of drone swarms in changing environments. To jointly manage these tasks, we propose a new synergetic optimization approach to coordinate drone swarms through Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a multi-agent DRL algorithm that builds upon on an iterative collective learning approach [7, 8]. It allows drones to determine long-term flying directions from departures to destinations (i.e., charging stations) via DRL, while autonomously evolving their short-term navigation and sensing policies based on shared plans from other drones through collective learning. The distributed sensing results generated during operations serve as rewards and observations for further DRL training. These results are aggregated and shared using a structured tree communication model, where each drone interacts with its parent and child nodes, making independent decisions. This ensures scalability (supporting a large number of agents), efficiency (low communication and computational cost), decentralization and resilience [9, 10, 11]. By employing this collaborative approach, collective learning reduces the frequency of decision-making and mitigates inefficient actions, effectively compensating for the limitations of DRL while retaining its strengths in long-term planning and optimization. In addition, a set of auxiliary approaches is introduced, including plan generation, plan selection, periodic state updates and a centralized training and decentralized execution module. To validate DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
The contributions of this paper are outlined as follows: (i) The first study of task assignment for large-scale, dynamic spatio-temporal sensing by a swarm of drones, concentrating on optimizing the flight paths, data collection strategies, and recharging. (ii) A novel synergetic optimization approach, DO-RL, that tackles this problem by integrating both multi-agent DRL and distributed optimization based on a structured tree communication model. (iii) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (iv) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (v) An open-source implementation of the proposed approach and an open dataset [12] containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
TABLE I: Comparison to related work: criteria covered (✓) or not (✕).
@@ II Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as the Traveling Salesmen Problem [1, 20]. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The evolutionary optimization methods, such as particle swarm optimization [21], and genetic algorithm [22], have been introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, events detection rate and the cost of flight paths. They rely on adaptive exploration of the solution space to optimize task allocation based on shared information and local interactions. However, these methods are centralized, wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms [23].
@@
Distributed optimization. Several distributed task assignment algorithms have been proposed. One such algorithm is the robust decentralized task assignment (RDTA) [13] that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another one is the consensus-based bundle algorithm (CBBA) [14], which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. Moreover, a planning-based coordination method using collective learning, named Economic Planning and Optimized Selections (EPOS), enables drones to optimize their navigation and sensing options under battery constraints [15, 24]. However, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often neglecting the need for strategic decision-making for overall benefits.
@@
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective for addressing distributed sensing with drones, managing complex optimization goals, energy constraints, and numerous heterogeneous agents. Ding et al. [16] integrated mobile crowdsensing (i.e., human equipped with mobile sensing devices) into UAV sensing, overcoming the limitations of both human mobility and drone battery life. Omoniwa et al. [25] presented a decentralized DRL approach that improves drone energy efficiency in providing wireless connectivity by sharing information with neighbouring drones. Zhao et al. [17] proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces. Their solution incorporates the task selection and scheduling using DRL, and adaptive recharging assignment. However, these approaches lack autonomy for drones in self-assignment tasks, reducing adaptability to the real-time environmental changes and increasing vulnerability to agent failures [23]. Additionally, they are confined to a small-scale task assignment problem due to the training complexity.
@@
Evolutionary algorithm and DRL. In this optimization process, the related synergetic optimization algorithms integrate the full learning and optimization capabilities of evolutionary algorithms and DRL. They independently solve sub-problems to generate partial solutions, which are then combined to form a complete solution [26]. Earlier work [27, 28] has proposed optimizing team policies using evolutionary algorithms with team rewards, while simultaneously optimizing individual policies through DRL with agent rewards, thus bypassing the non-stationary problem in multi-agent DRL. Li et al. [18] introduce a hybrid framework using shared observation encoders and independent linear policy representations to enhance collaboration on complex task, ensuring stable evolution through agent-level crossover and mutation. While benchmarks, such as MuJoCo [29] and Starcraft II [30], are useful for evaluation of these approaches, these benchmarks may struggle to adapt to the continuously evolving and unanticipated dynamics typical of real-world scenarios, such as fluctuating traffic patterns in drone-based monitoring systems.
@@
Transportation application. In the application of transportation system, drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors. Elloumi et al[31] proposed a road traffic monitoring system that generates adaptive drones trajectories by extracting information about vehicles. Samir et al. [19] leveraged a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. While these approaches demonstrate high performance in the detection of traffic vehicles, scalability becomes a challenge since they only deal with a small number of drones within a short duration.
@@ III-A Scenario and Assumptions
Matrix of required sensing values. In the context of a sensing task, each cell at a timeslot has specific sensing requirements that determine the data acquisition goal of drones. Such sensing requirements can be determined by city authorities as a continuous kernel density estimation, for example, monitoring cycling risk based on requirements calculated by past bike accident data and other information [15, 32]. The high risk level of a cell at a timeslot represents the high importance of sensing (e.g., the crucial intersection of traffic flow), and thus a high number of required sensing values is set. The matrix of required sensing values is denoted as $V(t)=\{V_{n,s}(t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$, where $V_{n,s}(t)$ denotes the required sensing value at cell $n$ and timeslot $s$ in the time period $t$. Based on actual sensing performance of drones according to Eq.(1), the sensing value collected by all drones observed by $u$ at cell $n$ and timeslot $s$ is formulated as:
@@
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a distinct altitude during its movement between cells to prevent mid-air collisions risks [33]. While previous work [34, 35] has explored a safe and cost-effective approach to a more realistic collision avoidance, this is not the primary focus of this paper. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. We assume that each charging period is of equal duration and provides sufficient time for the drones to be fully charged. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.
III-B Problem Formulation
@@
However, maximizing this metric leads to sensing imbalance and even blind areas. Some cells may be covered for a long time while some other cells may be never covered, i.e., the over-sensing and under-sensing [24]. Thus, it is necessary to consider matching indicators such as the sensing accuracy [24] to access such imbalances.
Sensing accuracy. It denotes the matching (correlation [^1] Error and correlation metrics (e.g. root mean squared error, cross-correlation or residuals of summed squares) estimate the matching, shown to be NP-hard combinatorial optimization problem in this context [7, 8].) between the total sensing values collected and the required ones. The metric is formulated as follows:
@@
Energy cost. It is the battery usage of drones to perform spatio-temporal sensing. A power consumption model [36] is used to calculate the power consumption with input the specification of drones (weight, propeller and power efficiency). This model estimates the cost of navigation and sensing plans, and emulates the outdoor environments [15]. The energy cost of each drone $u$ is formulated as:
@@ IV-A System Overview
IV-B Communication Model
To overcome these barriers, DO-RL leverages a structured tree communication model. Each drone, controlled by a local agent, locates and connects with other agents into a tree communication structure, as shown in Fig. 2. Once agents have established their proximity-based relationship, e.g., Euclidean distance, they are positioned starting from the leaves up to the root, each interacting with its children and parent in a bottom-up and top-down [11, 10]. For example, a drone $u$ aggregates and shares its plan $P^{u}$ in a bottom-up fashion, and then obtains the observed plan $P^{-u}$ in the top-down process [7]. These plans serve as the observation of an agent in DRL modeling. As this coordination among agents using hierarchical organization reduces the need for every node to communicate with every other node, the system is highly-efficient in collaboration and information exchange [7].
Through iterative bottom-up and top-down exchanges, agents can flexibly change their plans and adapt to dynamic sensing environments. This ensures that even if some nodes fail, the overall sensing performance remains effective. Moreover, the system can dynamically reconfigure, repositioning drones within the tree in response to communication failures, ensuring that a single drone failure has minimal impact on information flow. The resilience of this model has been previously demonstrated in [11] with large-scale real-world datasets, and thus this paper focuses on sensing performance and scalability of the proposed solution.
V Detailed System Design
V-A DRL Modeling
The core design of DO-RL lies in applying DRL on our problem. Once drones take actions of flying directions, they execute plans and return to charging stations, changing their current locations and battery levels while observing the navigation and sensing of other drones. Therefore, the problem scenario can be explained into a Markov decision process [6]. We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{u}(t)|_{u\in\mathcal{U}}\}$ is the current locations of drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station. $\mathcal{S}_{2}=\{e^{u}(t)|_{u\in\mathcal{U}}\}$ is the current battery levels of drones, which are calculated based on the energy cost. $\mathcal{S}_{3}=\{P^{u}(t)|_{u\in\mathcal{U}}\}$ is the plan of $u$. $\mathcal{S}_{4}=\{P^{-u}(t)|_{u\in\mathcal{U}}\}$ is the aggregated plan of other drones excluding $u$, which are shared via the structured communication model. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of period-by-period flying directions (N, E, S, W, NE, SE, SW, NW or O) determined by drones. (iii) Reward: Based on the objective function of Eq.(6), the expected immediate local reward of one drone at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\cdot\mathsf{Acc}(a^{u},t)-\alpha_{3}E(a^{u},t)$. Throughout the training process or episodes, the overall reward fluctuates based on the actions taken by the drones. This helps drones in prioritizing areas rich in sensor data by maximizing their reward, i.e., the highest cumulative overall performance.

### v7

@@ Abstract
Swarms of autonomous interactive drones can provide compelling sensing capabilities in Smart City applications, such as traffic monitoring. This paper focuses on the task assignment problem for large-scale spatio-temporal sensing by a drone swarm. However, existing approaches have distinct challenges: distributed evolutionary optimization, such as collective learning, lacks long-term adaptability in dynamic environments, while deep reinforcement learning (DRL) is limited to scale effectively due to the curse of dimensionality. Therefore, this paper proposes a novel synergetic optimization approach by integrating long-term DRL and short-term collective learning. Through this approach, each drone independently and proactively determines its flying direction and recharging location using DRL, while evolving their navigation and sensing policies through collective learning based on a structured tree communication model. Extensive experiments with datasets generated from realistic urban mobility demonstrate an outstanding performance of the proposed solution in complex scenarios. New insights show that this approach provides a win-win synthesis of short-term and long-term strategies for drone-based traffic monitoring, with short-term methods addressing training complexity and energy management, while long-term methods preserving high sensing performance.
keywords:
[…]
1 Introduction
@@
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as monitoring traffic congestion Barmpounakis and Geroliminis (2020), mapping natural disasters Deng et al. (2020) and smart agriculture Liu et al. (2024). Despite these advantages, drones are constrained by limited battery capacity, which impacts their spatio-temporal coverage. This limitation necessitates sophisticated autonomous control systems and precise task assignment within drone swarms. This involves the assignment of different sensing tasks to each drone while meeting the sensing requirements, drone capabilities and constraints Poudel and Moh (2022). Recently, scholars have solved the NP-hard combinatorial optimization problem of task assignment for spatio-temporal sensing, using evolutionary algorithms ranging from genetic algorithms to swarm intelligence methods Poudel and Moh (2022); Zhou et al. (2020). To enhance robustness and minimize the effects of individual drone failures, distributed optimization approaches have been introduced. These approaches allow drones to autonomously self-organize and self-assign tasks in a decentralized manner, preserving the autonomy of individual drones. However, these distributed approaches are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically.
@@
Strategic foresight in navigation and sensing is crucial for optimizing drone operations in dynamic environments, significantly influenced by flying directions and future sensor data requirements. For example, drones aware of an expected increase in traffic can proactively fly to those areas, improving tasks such as vehicle detection, even if the short-term benefits are not evident. This "slower is faster" effect emphasizes the importance of long-term planning for overall effectiveness in drone sensing. To address long-term optimization, deep reinforcement learning (DRL) has been introduced, utilizing the Bellman equation to consider cumulative rewards Cui et al. (2019). DRL based on deep neural networks support drones, as heterogeneous agents, to manage complex state spaces and adapt to changing environments.
@@
Therefore, in this paper, we address a novel task assignment problem for large-scale spatio-temporal sensing, focusing on the navigation (determining flight path), sensing (scheduling data collection) and recharging (choosing charging stations as destinations) of drone swarms in changing environments. To jointly manage these tasks, we propose a new synergetic optimization approach to coordinate drone swarms through Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a multi-agent DRL algorithm that builds upon on an iterative collective learning approach Pournaras et al. (2018); Pournaras (2020). It allows drones to determine long-term flying directions from departures to destinations (i.e., charging stations) via DRL, while autonomously evolving their short-term navigation and sensing policies based on shared plans from other drones through collective learning. The distributed sensing results generated during operations serve as rewards and observations for further DRL training. These results are aggregated and shared using a structured tree communication model, where each drone interacts with its parent and child nodes, making independent decisions. This ensures scalability (supporting a large number of agents), efficiency (low communication and computational cost), decentralization and resilience Majumdar et al. (2023); Pournaras et al. (2020); Nikolic and Pournaras (2019). By employing this collaborative approach, collective learning reduces the frequency of decision-making and mitigates inefficient actions, effectively compensating for the limitations of DRL while retaining its strengths in long-term planning and optimization. In addition, a set of auxiliary approaches is introduced, including plan generation, plan selection, periodic state updates and a centralized training and decentralized execution module. To validate DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
The contributions of this paper are outlined as follows: (i) The first study of task assignment for large-scale, dynamic spatio-temporal sensing by a swarm of drones, concentrating on optimizing the flight paths, data collection strategies, and recharging. (ii) A novel synergetic optimization approach, DO-RL, that tackles this problem by integrating both multi-agent DRL and distributed optimization based on a structured tree communication model. (iii) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (iv) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (v) An open-source implementation of the proposed approach and an open dataset [^1] https://doi.org/10.6084/m9.figshare.24476533.v2 containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
Table 1: Comparison to related work: criteria covered (✓) or not (✕).
@@ 2 Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as the Traveling Salesmen Problem Bartolini et al. (2019). This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The evolutionary optimization methods, such as particle swarm optimization Phung and Ha (2021), and genetic algorithm Pehlivanoglu and Pehlivanoglu (2021), have been introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, events detection rate and the cost of flight paths. They rely on adaptive exploration of the solution space to optimize task allocation based on shared information and local interactions. However, these methods are centralized, wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms Omidshafiei et al. (2017).
@@
Distributed optimization. Several distributed task assignment algorithms have been proposed. One such algorithm is the robust decentralized task assignment (RDTA) Alighanbari and How (2006) that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another one is the consensus-based bundle algorithm (CBBA) Chen et al. (2022), which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. Moreover, a planning-based coordination method using collective learning, named Economic Planning and Optimized Selections (EPOS), enables drones to optimize their navigation and sensing options under battery constraints Qin et al. (2022); Qin and Pournaras (2023). However, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often neglecting the need for strategic decision-making for overall benefits.
@@
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective for addressing distributed sensing with drones, managing complex optimization goals, energy constraints, and numerous heterogeneous agents. Ding et al. Ding et al. (2021) integrated mobile crowdsensing (i.e., human equipped with mobile sensing devices) into UAV sensing, overcoming the limitations of both human mobility and drone battery life. Omoniwa et al. Omoniwa et al. (2023) presented a decentralized DRL approach that improves drone energy efficiency in providing wireless connectivity by sharing information with neighbouring drones. Zhao et al. Zhao et al. (2022) proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces. Their solution incorporates the task selection and scheduling using DRL, and adaptive recharging assignment. However, these approaches lack autonomy for drones in self-assignment tasks, reducing adaptability to the real-time environmental changes and increasing vulnerability to agent failures Omidshafiei et al. (2017). Additionally, they are confined to a small-scale task assignment problem due to the training complexity.
@@
Evolutionary algorithm and DRL. In this optimization process, the related synergetic optimization algorithms integrate the full learning and optimization capabilities of evolutionary algorithms and DRL. They independently solve sub-problems to generate partial solutions, which are then combined to form a complete solution Li et al. (2024). Earlier work Majumdar et al. (2020); Aydeniz et al. (2023) has proposed optimizing team policies using evolutionary algorithms with team rewards, while simultaneously optimizing individual policies through DRL with agent rewards, thus bypassing the non-stationary problem in multi-agent DRL. Li et al. Li et al. (2023) introduce a hybrid framework using shared observation encoders and independent linear policy representations to enhance collaboration on complex task, ensuring stable evolution through agent-level crossover and mutation. While benchmarks, such as MuJoCo Peng et al. (2021) and Starcraft II Samvelyan et al. (2019), are useful for evaluation of these approaches, these benchmarks may struggle to adapt to the continuously evolving and unanticipated dynamics typical of real-world scenarios, such as fluctuating traffic patterns in drone-based monitoring systems.
@@
Transportation application. In the application of transportation system, drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors. Bakirci et alBakirci (2024) explored vehicle detection using the YOLOv8 algorithm on aerial images captured by a custom UAV. Samir et al. Samir et al. (2020) leveraged a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. While these approaches demonstrate high performance in the detection of traffic vehicles, scalability becomes a challenge since they only deal with a small number of drones within a short duration.
@@ 3.1 Scenario and Assumptions
Matrix of required sensing values. In the context of a sensing task, each cell at a timeslot has specific sensing requirements that determine the data acquisition goal of drones. Such sensing requirements can be determined by city authorities as a continuous kernel density estimation, for example, monitoring cycling risk based on requirements calculated by past bike accident data and other information Castells-Graells et al. (2020). The high risk level of a cell at a timeslot represents the high importance of sensing (e.g., the crucial intersection of traffic flow), and thus a high number of required sensing values is set. The matrix of required sensing values is denoted as $V(t)=\{V_{n,s}(t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$, where $V_{n,s}(t)$ denotes the required sensing value at cell $n$ and timeslot $s$ in the time period $t$. Based on actual sensing performance of drones according to Eq.(1), the sensing value collected by all drones observed by $u$ at cell $n$ and timeslot $s$ is formulated as:
@@
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a distinct altitude during its movement between cells to prevent mid-air collisions risks Tang et al. (2021). While previous work Qin et al. (2024); Batinovic et al. (2022) has explored a safe and cost-effective approach to a more realistic collision avoidance, this is not the primary focus of this paper. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. We assume that each charging period is of equal duration and provides sufficient time for the drones to be fully charged. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.
3.2 Problem Formulation
@@
However, maximizing this metric leads to sensing imbalance and even blind areas. Some cells may be covered for a long time while some other cells may be never covered, i.e., the over-sensing and under-sensing Qin and Pournaras (2023). Thus, it is necessary to consider matching indicators such as the sensing accuracy Qin and Pournaras (2023) to access such imbalances.
Sensing accuracy. It denotes the matching (correlation [^2] Error and correlation metrics (e.g. root mean squared error, cross-correlation or residuals of summed squares) estimate the matching, shown to be NP-hard combinatorial optimization problem in this context Pournaras et al. (2018); Pournaras (2020).) between the total sensing values collected and the required ones. The metric is formulated as follows:
@@
Energy cost. It is the battery usage of drones to perform spatio-temporal sensing. A power consumption model Monwar et al. (2018) is used to calculate the power consumption with input the specification of drones (weight, propeller and power efficiency). This model estimates the cost of navigation and sensing plans, and emulates the outdoor environments Qin et al. (2022). The energy cost of each drone $u$ is formulated as:
@@ 4.1 System Overview
4.2 Communication Model
To overcome these barriers, DO-RL leverages a structured tree communication model. Each drone, controlled by a local agent, locates and connects with other agents into a tree communication structure, as shown in Fig. 2. Once agents have established their proximity-based relationship, e.g., Euclidean distance, they are positioned starting from the leaves up to the root, each interacting with its children and parent in a bottom-up and top-down Nikolic and Pournaras (2019); Pournaras et al. (2020). For example, a drone $u$ aggregates and shares its plan $P^{u}$ in a bottom-up fashion, and then obtains the observed plan $P^{-u}$ in the top-down process Pournaras et al. (2018). These plans serve as the observation of an agent in DRL modeling. As this coordination among agents using hierarchical organization reduces the need for every node to communicate with every other node, the system is highly-efficient in collaboration and information exchange Pournaras et al. (2018).
Through iterative bottom-up and top-down exchanges, agents can flexibly change their plans and adapt to dynamic sensing environments. This ensures that even if some nodes fail, the overall sensing performance remains effective. Moreover, the system can dynamically reconfigure, repositioning drones within the tree in response to communication failures, ensuring that a single drone failure has minimal impact on information flow. The resilience of this model has been previously demonstrated in Nikolic and Pournaras (2019) with large-scale real-world datasets, and thus this paper focuses on sensing performance and scalability of the proposed solution.
@@ 4.2 Communication Model
5 Detailed System Design
5.1 DRL Modeling
The core design of DO-RL lies in applying DRL on our problem. Once drones take actions of flying directions, they execute plans and return to charging stations, changing their current locations and battery levels while observing the navigation and sensing of other drones. Therefore, the problem scenario can be explained into a Markov decision process Cui et al. (2019). We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{u}(t)|_{u\in\mathcal{U}}\}$ is the current locations of drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station. $\mathcal{S}_{2}=\{e^{u}(t)|_{u\in\mathcal{U}}\}$ is the current battery levels of drones, which are calculated based on the energy cost. $\mathcal{S}_{3}=\{P^{u}(t)|_{u\in\mathcal{U}}\}$ is the plan of $u$. $\mathcal{S}_{4}=\{P^{-u}(t)|_{u\in\mathcal{U}}\}$ is the aggregated plan of other drones excluding $u$, which are shared via the structured communication model. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of period-by-period flying directions (N, E, S, W, NE, SE, SW, NW or O) determined by drones. (iii) Reward: Based on the objective function of Eq.(6), the expected immediate local reward of one drone at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\alpha_{1}\mathsf{Eff}(a^{u},t)+\alpha_{2}\cdot\mathsf{Acc}(a^{u},t)-\alpha_{3}E(a^{u},t)$. Throughout the training process or episodes, the overall reward fluctuates based on the actions taken by the drones. This helps drones in prioritizing areas rich in sensor data by maximizing their reward, i.e., the highest cumulative overall performance.